OpenStudy (anonymous):

PLEASE HELP!!!!!!!! In a water treatment plant, water passes through a cone-shaped filter with a height of 4 m and a diameter of 9 m. The water flows from the filter into a cylinder below it. One full cone-shaped filter fills one-fifth of the cylinder. What are the dimensions of the cylinder? Use 3.14 to approximate pi. A. h = 3 m; r = 3 m B. h = 4.23 m; r = 5 m C. h = 15 m; r = 6 m D. h = 33.75 m; r = 2 m

OpenStudy (triciaal):

volume of a cylinder = pi*r^2*h volume of a cone = 1/3*pi*r^2*h volume of cone = 1/3*volume of cylinder

OpenStudy (anonymous):

but i still dont know what it is

OpenStudy (triciaal):

find volume of cone from numbers given use ratio volume of this cone = 1/5 the volume of this cylinder the diameter is twice the radius

OpenStudy (anonymous):

because this question makes no sense

OpenStudy (triciaal):

I think the radius of the cone and the cylinder should be the same but 9/2 = 4.5 not 5 unless they approximate but on second thought it should be smaller

OpenStudy (anonymous):

@triciaal can u help me?

OpenStudy (triciaal):